OpenStudy (anonymous):

Simplify sqrrt75x^8y^11 Assume that all variables represent positive real numbers.

OpenStudy (anonymous):

ahh one of these again :P gotta love them haaha ok here is what we got: \[\sqrt{75x^8y ^{11}}\] The part that is tedious about this is that you gotta split the variables up into seprate squares with radicals if you were too youd end up with 4 x^2 radicals and 5 y^2 radicals the 75 can have a radical 25 and radical 3 taken out where the radical 25 simplifies to 5 so this is what we have: \[5\sqrt{3} * x * x * x * x * y * y * y * y * y * \sqrt{y}\] Combine everything together to get \[5x^4y^5\sqrt{3y}\]

OpenStudy (anonymous):

Thanks very much

OpenStudy (anonymous):

Welcome forgot to explain the radical y is sitll there due to the power of 11 originally being odd so.